# Uncertainty Quantification for Evaluating Machine Translation Bias

### v1

@@ Abstract
In machine translation (MT), when the source sentence includes a lexeme whose gender is not overtly marked, but whose target-language equivalent requires gender specification, the model must infer the appropriate gender from the context and/or external knowledge. Studies have shown that MT models exhibit biased behaviour, relying on stereotypes even when they clash with contextual information. We posit that apart from confidently translating using the correct gender when it is evident from the input, models should also maintain uncertainty about the gender when it is ambiguous. Using recently proposed metrics of semantic uncertainty, we find that models with high translation and gender accuracy on unambiguous instances do not necessarily exhibit the expected level of uncertainty in ambiguous ones. Similarly, debiasing has independent effects on ambiguous and unambiguous translation instances. [^1] The code will be available at https://anonymous.4open.science/r/uncertainty_bias_ambiguity-8A4C/
Uncertainty Quantification for Evaluating Machine Translation Bias
@@ 1 Introduction
Language is inherently ambiguous, and meaning is often resolved through context. However, not all ambiguity is resolvable (van Deemter, 1998). When humans process language, they draw on linguistic, cognitive, and social biases to arrive at an interpretation (Cairns, 1973). While linguistic biases ease cognitive processing, some can also have harmful effects, such as reinforcing existing social inequalities (Beukeboom, 2013). NLP models exhibit sensitivity to many of the human biases (Echterhoff et al., 2024), and even exaggerate them (Dhamala et al., 2021), as well as introduce additional ones (Tjuatja et al., 2024). When the input is unresolvably ambiguous, favoring a single output necessarily relies on biases. Therefore, a well-designed model should refrain from making a single prediction, instead requesting clarification or generating multiple alternative outputs.
Most studies on decoding with language models (LMs) for machine translation (MT) evaluate a single prediction, usually generated with beam search, per translation instance. Consequently, most studies on uncertainty quantification (UQ) use uncertainty to predict the quality of the translation recovered by beam search (Fomicheva et al., 2020; Cheng and Vlachos, 2024). Previous work on bias in MT has focused on LM performance against gold standard labels (Stanovsky et al., 2019), and previous work on ambiguity in MT has likewise focused on resolvable cases (Barua et al., 2024; Martelli et al., 2025). Thus, most work on uncertainty and ambiguity assumes that there is a single correct translation per instance. However, less attention has been given to ambiguous source sentences, where the choice of the LM cannot be guaranteed to be correct without additional context. This may be seen as a form of aleatoric uncertainty (Hora, 1996), uncertainty which is inherent in the data and irreducible. According to Baan et al. (2024), the spread of probability mass in LMs represents both lack of confidence as well as variation in human generations. We adopt this line of thinking and examine whether LMs accurately and fairly represent the range of possible translations for ambiguous source sentences.
@@
In this work, we leverage distribution-level uncertainty metrics to evaluate cases where the model should not be certain about its predictions due to ambiguity in the input. Figure 1 shows two versions of a sentence, with unambiguous (top) or ambiguous (bottom) gender of the noun ‘protester’, and the different probabilities assigned to the Spanish translations of the determiner of this noun. An ideal model should assign a higher probability for the feminine determiner (‘la’) when the gender is disambiguated by the pronoun, and produce equal probabilities for masculine and feminine translations when the gender is ambiguous. However, state-of-the-art MT models, including debiased ones, tend to produce more uniform probability distributions for unambiguous inputs, and less uniform distributions for ambiguous ones. This indicates that the model probabilities are influenced by stereotypical associations between protesting and masculinity, causing the models to default to the masculine form even when no gender preference is warranted, and to select the feminine form with low confidence despite clear contextual cues.
@@
To study this systematically, we focus on translating sentences from a language that does not mark gender in nouns and verbs (English) into languages that do (Spanish, French, Ukrainian, and Russian). We use the WinoMT dataset (Stanovsky et al., 2019), which includes stereotypical gender roles, and extend it with manual translations and automatic annotations of additional cognitive bias cues, such as implicit causality verbs. In some cases, the gender is resolvable from context, while in others it is not. We explore different variants of Semantic Uncertainty metrics (Cheng and Vlachos, 2024; Farquhar et al., 2024) to quantify the semantic diversity of translation samples, finding that these metrics effectively capture the variation in gender caused by bias triggers. We validate the metrics against the established gender accuracy metric and account for the effects of translation accuracy. Our main findings are: 1) stereotypes and linguistic biases influence gender translation, 2) the degree of bias corresponds to overall model translation accuracy in unambiguous cases, 3) the degree of bias corresponds to translation accuracy at the instance level in ambiguous cases, 4) debiasing effects vary depending on input ambiguity, translation accuracy, and target language.
@@ 2 Related Work
Researchers have addressed various biases in MT, including algorithmic bias (Vanmassenhove et al., 2021), as well as gender, number and formality biases (Měchura, 2022). Gender stereotypes are triggered not only by semantic content, but also by speech mannerisms (Dawkins et al., 2024) and person names (Saunders and Olsen, 2023). Existing solutions for gender translation in both unambiguous (Robinson et al., 2024) and ambiguous cases (Cho et al., 2019; Gonen and Webster, 2020; Vanmassenhove and Monti, 2021) rely on tools or human annotation of gender, limiting their generalisability to other types of ambiguities. For example, Cho et al. (2019) explore unresolvable ambiguity by generating multiple translations for a sentence with an ambiguous pronoun, however, it is restricted to specific sentence types.
@@
In NLP and ML research, methods for distinguishing aleatoric and epistemic uncertainty have been proposed (Hou et al., 2024), however they do not distinguish between data randomness and data ambiguity. Some work has made the link between biases and uncertainty (Sicilia et al., 2024; Kuzucu et al., 2025), as well as between ambiguity and uncertainty (Kim, 2025; Cheng and Amiri, 2024), however, these papers construe uncertainty as a signal for poor performance and ambiguity as low quality inputs, which differs from our definition of ambiguity as an indispensable feature of language. Work in Question Answering has found that the best methods for detecting ambiguous inputs involve quantifying repetition within sampled model outputs (Cole et al., 2023), and using white-box metrics such as entropy (Yang et al., 2025).
@@
UQ in MT has been used as a proxy for Quality Estimation (QE). For example, Fomicheva et al. (2020) use MT model uncertainty to estimate translation quality without references, while Glushkova et al. (2021) apply the same technique to the uncertainty of the QE models themselves. Other approaches use UQ to identify difficult instances and enhance training by applying curriculum learning (Zhou et al., 2020), semantic augmentation (Wei et al., 2020), balancing of multilingual training data (Wu et al., 2021) or test-time adaptation (Zhan et al., 2023). Wang et al. (2024b) examine zero-shot translation and distinguish between model uncertainty and data uncertainty, however their focus with regard to data uncertainty is on noisy, low-quality training data rather than inherent ambiguity. Cognitive science research has demonstrated that entropy-based uncertainty metrics are a suitable measure of ambiguity in human translations Bangalore et al. (2016), but this insight has yet to be applied in MT. To the best of our knowledge, no prior UQ-based approach in MT has explored ambiguity as a particular type of data uncertainty.
@@ 3 Method
We propose to quantify gender bias in Neural Machine Translation (NMT) models by characterising how gender is assigned to nouns across the predictive distribution. To do so, we base our methods on recently proposed UQ metrics which are founded on the classic Shannon entropy but take into account similarities between random Monte Carlo samples from the model. We first provide a brief overview of these UQ methods.
@@
$\mathcal{H}(\mathcal{Y})=\underset{y\sim\mathcal{Y}}{\mathbb{E}}\left[I(y)\right],$
@@
Semantic Entropy (se; Farquhar et al., 2024) identifies semantic equivalences between elements and clusters them together according to a textual entailment model, mapping each $y$ to a cluster $c$. In our implementation we use a multilingual mDeberta model (He et al., 2021) finetuned on the Natural Language Inference (NLI) task by Laurer et al. (2022). Then, surprisal is the negative log probability of an element being in $c$:
$I_{\mathrm{SE}}(y)=-\log\underset{y^{\prime}\sim\mathcal{Y}}{\mathbb{E}}1\left[y^{\prime}\in{c}\right].$
Similarity-sensitive Shannon Entropy (s3e; Ricotta and Szeidl, 2006; Cheng and Vlachos, 2024) sets the surprisal of $y$ to the negative log of its expected similarity with all other outputs:
$I_{\mathrm{S3E}}(y)=-\log\mathbb{E}_{y^{\prime}\sim\mathcal{Y}}\left[\mathcal{S}(y,y^{\prime})\right],$
@@
where $\mathcal{S}$ is a similarity function satisfying $\mathcal{S}(y,y^{\prime})\in[0,1]$ and $\mathcal{S}(y,y^{\prime})=1$ if $y=y^{\prime}$. Following Cheng and Vlachos (2024), we use cosine similarity between sentence embeddings of $y$ and $y^{\prime}$ generated by a multilingual E5 text embedding model (Wang et al., 2024a).
@@
We also define Gender Entropy (ge), which is calculated like se but clusters elements based on the gender class of the translated focus noun. To determine the gender class, we use Spacy [^2] https://spacy.io/ and pymorphy2 (Korobov, 2015) morphological parsers.
@@
se, s3e, and ge must be approximated with random sampling from $p(\cdot|x)$, which we perform with $\epsilon$-sampling (Hewitt et al., 2022), drawing 128 samples per source sentence. Further details about these UQ methods are given in Appendix A.
@@
Our gender bias metrics are based on surprisal and entropy given by these UQ methods. The first desideratum is that for source sentences with unambiguous gender, an unbiased model should have lower surprisal of a translation with correct gender inflection compared to an incorrect inflection. Therefore, unbiased models should minimise relative surprisal, defined as:
$\Delta I=\frac{I(y_{\text{correct}})-I({y_{\text{incorrect}}})}{\frac{1}{2}(I({y_{\text{correct}})}+I({y_{\text{incorrect}}})}.$
The second desideratum is that the entropy of unbiased models should not be affected by the presence of bias cues. Thus, we define normalised entropy, which compares the $\mathcal{H}$ of a source sentence $x$ to the average entropy across its contrast set $\mathcal{G}_{x}$. $\mathcal{G}_{x}$ is a group of minimally different sentences that are identical to $x$ apart from the pronoun (e.g., ‘she’, ‘he’, ‘they’), including $x$ itself. The three sentences in Table 1 comprise a $\mathcal{G}_{x}$. Formally:
[…]
@@
The third desideratum is that models should show higher uncertainty for an input which is ambiguous with regard to gender, as compared to an input which is unambiguous, disregarding all biases in the input. It should therefore minimise relative entropy, defined as:
@@
$\Delta\mathcal{H}=\frac{\mathcal{H}(\mathcal{Y}_{\text{unambiguous}})-\mathcal{H}(\mathcal{Y}_{\text{ambiguous}})}{\frac{1}{2}\left(\mathcal{H}(\mathcal{Y}_{\text{unambiguous}})+\mathcal{H}\left(\mathcal{Y}_{\text{ambiguous}}\right)\right)}.$

### v2

# Uncertainty Quantification for Evaluating Gender Bias in Machine Translation

@@ Abstract
The predictive uncertainty of machine translation (MT) models is typically used as a quality estimation proxy. In this work, we posit that apart from confidently translating when a single correct translation exists, models should also maintain uncertainty when the input is ambiguous. We use uncertainty to measure gender bias in MT systems. When the source sentence includes a lexeme whose gender is not overtly marked, but whose target-language equivalent requires gender specification, the model must infer the appropriate gender from the context and can be susceptible to biases. Prior work measured bias via gender accuracy, however it cannot be applied to ambiguous cases. Using semantic uncertainty, we are able to assess bias when translating both ambiguous and unambiguous source sentences, and find that high translation accuracy does not correlate with exhibiting uncertainty appropriately, and that debiasing affects the two cases differently. [^1] The code will be available at https://anonymous.4open.science/r/uncertainty_bias_ambiguity-8A4C/
Uncertainty Quantification for Evaluating Gender Bias
in Machine Translation
@@ 1 Introduction
Language is inherently ambiguous, and meaning is often resolved through context. However, not all ambiguity is resolvable (van Deemter, 1998). Preferring one interpretation of an ambiguous statement over another relies on biases (Cairns, 1973). Natural Language Processing (NLP) models have been shown not only to reproduce existing social inequalities (Beukeboom, 2013) but also to amplify them (Dhamala et al., 2021; Tjuatja et al., 2024). Therefore, when faced with an ambiguous input, a well-designed model should remain uncertain to avoid reinforcing biases.
Uncertainty that arises from unresolvable ambiguity can be understood as a form of aleatoric uncertainty (Hora, 1996), since it is inherent to the data and cannot be reduced. Most research considers model uncertainty to be a proxy for prediction correctness (Kumar et al., 2019), although work has shown that model probability encodes both aleatoric and epistemic uncertainty (Baan et al., 2024).
Machine Translation (MT) provides an ideal testing ground for uncertainty caused by ambiguity. That is because languages differ in the number of grammatical categories they explicitly mark, and translating into a richer language requires disambiguation. This allows for explicitly measuring uncertainty caused by ambiguity by changing the specificity of grammatical gender and comparing the resulting translations.
@@
In this work, we leverage uncertainty quantification (UQ) to evaluate bias in cases where the model ought not to be certain about its predictions due to ambiguity in the input. Figure 1 shows two versions of a sentence, with unambiguous (top) or ambiguous (bottom) gender of the noun ‘protester’, and the different probabilities assigned to the Spanish translations of the determiner of this noun by an Opus-MT translation model (Tiedemann and Thottingal, 2020). An ideal model should assign a higher probability for the feminine determiner (‘la’) when the gender is disambiguated by the pronoun and produce equal probabilities for masculine and feminine translations when the gender is ambiguous. However, in our preliminary analyses, we observed that state-of-the-art MT models, including debiased variants, tend to produce more uniform probability distributions for unambiguous inputs, and less uniform distributions for ambiguous ones. This indicates that model probabilities are influenced by stereotypical associations between protesting and masculinity, causing the models to default to the masculine form even when no gender preference is warranted (bottom), and predict the feminine form with relatively low confidence despite very clear contextual cues (top).
In our analysis we consider semantic UQ metrics for a number of reasons. First, entropy-based uncertainty metrics have been shown to be a suitable measure of ambiguity in human translations (Bangalore et al., 2016), however it has not been applied in MT. Second, quantifying repetition within sampled model outputs, as well as using white-box metrics such as entropy, has been successfully applied to detecting ambiguous inputs in question answering (Cole et al., 2023; Yang et al., 2025). Third, semantic UQ metrics have been shown to capture linguistic nuance in generation tasks, where the diversity of samples is better estimated if their semantics are taken into account (Cheng and Vlachos, 2024; Farquhar et al., 2024).
Our goal to detect bias using semantic UQ metrics. We focus on extrinsic bias (Goldfarb-Tarrant et al., 2021), which manifests in model outputs that are affected by stereotypes. The currently established metric of gender accuracy evaluates gender bias by measuring how often the model generates the correct gender in the translation (Stanovsky et al., 2019). This metric can only be applied in cases where there exists a correct gender, i.e. unambiguous instances. Bias also appears in ambiguous scenarios, therefore we propose to measure it across the board.
@@
To study this systematically, we focus on translating sentences from a language that does not mark gender in nouns and verbs (English) into languages that do (Spanish, French, Ukrainian, Russian, German, Italian and Greek). We use the WinoMT (Stanovsky et al., 2019) as well as the mGeNTe (Savoldi et al., 2025) datasets. Both datasets contain instances where the gender is resolvable from context as well as instances where it is not. mGeNTe provides alternative translations with gender either explicitly expressed in the target sentence or avoided in the translation. WinoMT includes stereotypical gender roles annotations, and we additionally extend it with manual translations. We first validate the usefulness of semantic UQ metrics by showing that they correlate well with gender accuracy on unambiguous items across different models. Then, we use these metrics to investigate the biases exhibited by the models in ambiguous cases. Our main findings are: 1) the degree of bias correlates with overall model translation accuracy in unambiguous cases, 2) the degree of bias corresponds to translation accuracy at the instance level in ambiguous cases, 3) debiasing effects vary depending on input ambiguity, translation accuracy, and target language.
@@ 2 Related Work
Bias in MT has been classified into under-representation and stereotyping, which is persistent even in resolvable settings (Savoldi et al., 2021; Ghosh and Caliskan, 2023). Various methods for debiasing have been proposed (Saunders and Byrne, 2020), however the evaluation of bias is not uniform. Earlier work in gender bias in MT has used intrinsic bias evaluation by testing the embeddings of gendered terms and their association with certain topics (Ramesh et al., 2021), and extrinsic bias evaluation by counting correct translations against a target (Wang et al., 2022). Gender ambiguity in MT has mostly been addressed where it is resolvable (Currey et al., 2022; Robinson et al., 2024), similarly to other types of lexical ambiguity (Barua et al., 2024; Martelli et al., 2025). Existing solutions for unresolvable ambiguous cases (Cho et al., 2019; Gonen and Webster, 2020; Vanmassenhove and Monti, 2021; Farkas and Németh, 2022; Piazzolla et al., 2023) rely on tools or human annotation of gender on specific words, limiting their generalisability to other types of ambiguities. In contrast, some work circumvents the problem of unresolvable gender ambiguity. Lauscher et al. (2023) translate the neutral pronoun while preserving the neutrality in the target language, avoiding exhibiting the bias altogether, while Alhafni et al. (2022) rewrite gender according to the preferences of the user and Vanmassenhove et al. (2018) include the speaker gender information. Stafanovičs et al. (2020) propose a solution for unresolvable ambiguity by decoupling the task of determining the gender information present in the input from translation, this way circumventing the issue of what should be done when gender information is not present.
@@
When it comes to uncertainty in ML models, methods for distinguishing aleatoric and epistemic uncertainty have been proposed (Hou et al., 2024), however they do not distinguish between data randomness and data ambiguity. Some work has made the link between biases and uncertainty by showing that model uncertainty is linked to the sycophancy bias (Sicilia et al., 2025) as well as the fact that model uncertainty differs when the input mentions different demographic groups (Kuzucu et al., 2025). Ambiguity has also been shown to be reflected in uncertainty in different NLP tasks (Kim, 2025; Cheng and Amiri, 2024), however, these studies construe uncertainty as a signal for ambiguity stemming from low quality data, not including the ambiguity which is an indispensable feature of language.
@@
UQ in MT has been used as a proxy for Quality Estimation (QE). For example, Fomicheva et al. (2020) use MT model uncertainty to estimate translation quality without references, while Glushkova et al. (2021) apply the same technique to the uncertainty of the QE models themselves. Other approaches use UQ to identify difficult instances and enhance training (Zhou et al., 2020; Wei et al., 2020; Wu et al., 2021; Zhan et al., 2023). Wang et al. (2024b) examine zero-shot translation and distinguish between model uncertainty and data uncertainty, however their focus with regard to data uncertainty is on noisy, low-quality training data rather than inherent ambiguity. Ott et al. (2018) show that uncertainty caused by low quality training data is linked to a deterioration of accuracy in beam search with large beams. To the best of our knowledge, no prior UQ-based approach in MT has explored ambiguity as a particular type of data uncertainty.
@@ 3 Method
We propose to quantify gender bias in Neural Machine Translation (NMT) models by characterising how gender is assigned to nouns across the predictive distribution. Our bias evaluation method rests on two conditions:
Condition 1: For source sentences with unambiguous gender, an unbiased model should prefer a translation with correct gender inflection compared to an incorrect inflection.
Condition 2: Unbiased models should show higher uncertainty for inputs with ambiguous gender, as compared to unambiguous ones, disregarding the biases in the input.
To do so, we base our approach on recently proposed UQ metrics which are founded on the classic Shannon entropy but take into account similarities between random Monte Carlo samples from the model. Sampling-based metrics are preferred over deterministic decoding methods, as they better capture ambiguity in the ground truth and leverage model stochasticity. We first provide a brief overview of the UQ methods.
@@
$\mathcal{H}(\mathcal{Y})=\underset{y\sim\mathcal{Y}}{\mathbb{E}}\left[I(y)\right],$ (1)
@@
Semantic Entropy (se) (Farquhar et al., 2024) is a metric designed to quantify uncertainty by identifying semantic equivalences among output elements. It achieves this by grouping elements $y$ into clusters $c$ of semantically equivalent outputs based on a textual entailment model. To implement this clustering, we use a multilingual mDeberta model (He et al., 2021) finetuned on the Natural Language Inference (NLI) task (Laurer et al., 2022). se is intended to indirectly capture clusters of translations based on the gender of the subject or object, given that two identical translations that only differ by gender marking would not entail one another.
Given the cluster assignment, the surprisal of an output $y$ under Semantic Entropy is defined as the negative logarithm of the expected probability mass assigned to the cluster $c$ to which $y$ belongs:
$I_{\mathrm{SE}}(y)=-\log\mathbb{E}_{y^{\prime}\sim\mathcal{Y}}\left[1[y^{\prime}\in c]\right],$ (2)
where $\mathcal{Y}$ is the predictive distribution over outputs and $1[\cdot]$ is the indicator function.
Complementing se, Similarity-sensitive Shannon Entropy (s3e) (Ricotta and Szeidl, 2006; Cheng and Vlachos, 2024) generalizes surprisal by incorporating graded semantic similarity between outputs rather than binary clusters, capturing subtler translation differences such as gender marking on verbs and adjectives. Specifically, the surprisal of $y$ is the negative logarithm of the expected similarity between $y$ and other outputs $y^{\prime}$:
$I_{\mathrm{S3E}}(y)=-\log\mathbb{E}_{y^{\prime}\sim\mathcal{Y}}\left[\mathcal{S}(y,y^{\prime})\right],$ (3)
@@
where the similarity function $\mathcal{S}:\mathcal{Y}\times\mathcal{Y}\to[0,1]$ satisfies $\mathcal{S}(y,y^{\prime})=1$ if and only if $y=y^{\prime}$. Following Cheng and Vlachos (2024), we compute $\mathcal{S}$ as the cosine similarity between sentence embeddings of $y$ and $y^{\prime}$, generated by a multilingual E5 text embedding model (Wang et al., 2024a).
@@
We also define Gender Entropy (ge), which is calculated like se but clusters elements based on the gender class of the translated noun. To determine the gender class, we use Spacy [^2] https://spacy.io/ and pymorphy2 (Korobov, 2015) morphological parsers. ge provides the most direct way to measure uncertainty with regard to the gender.
@@
To calculate se, s3e and ge we approximate the model’s predictive distribution $p(\cdot\mid x)$ by drawing $N=128$ samples per input using stochastic decoding. Specifically, we perform $\epsilon$-sampling (Hewitt et al., 2022) with top-$p$ sampling (epsilon_cutoff = 0.2). Further details about these UQ methods are given in Appendix B.
We also compute the log probability of each generated sequence by summing the log probabilities of its tokens. Formally, given a sequence $s=(w_{1},\ldots,w_{T})$ and model output logits $\text{scores}_{t}$ at step $t$, we calculate
$\log p(s)=\sum_{t=1}^{T}\log p(w_{t}\mid w_{<t},x),$
where $\log p(w_{t}\mid w_{<t},x)$ is obtained by applying the log-softmax to $\text{scores}_{t}$. This provides a direct likelihood measure of the sequence, used as a simpler alternative to sampling-based surprisal and entropy metrics.
@@
Our gender bias metrics are based on surprisal and entropy from these UQ methods. Thus, to meet Condition 1, in an ambiguous setting, an unbiased model should assign lower surprisal to a translation with the correct gender than to an incorrect one. Accordingly, unbiased models should minimize relative surprisal, defined as:
$\Delta I=\frac{I(y_{\text{correct}})-I(y_{\text{incorrect}})}{\frac{1}{2}\left(I(y_{\text{correct}})+I(y_{\text{incorrect}})\right)}.$ (4)
@@
To meet the Condition 2, unbiased models should show higher uncertainty for an input which is ambiguous with regard to gender, as compared to an input which is unambiguous. It should therefore minimise relative entropy, defined as:
@@
$\Delta\mathcal{H}=\frac{\mathcal{H}(\mathcal{Y}_{\text{unambiguous}})-\mathcal{H}(\mathcal{Y}_{\text{ambiguous}})}{\frac{1}{2}\left(\mathcal{H}(\mathcal{Y}_{\text{unambiguous}})+\mathcal{H}\left(\mathcal{Y}_{\text{ambiguous}}\right)\right)}.$ (5)